# Strand, London WC2R 2LS, UK

### v1

@@ Acknowledgements
It is a pleasure to thank Michael Anderson, Dionysios Anninos, Pau Figueras, Davide Gaiotto, Chawakorn Maneerat, Edgar Shaghoulian and Eva Silverstein for useful discussions. TZ especially thanks Sergio Sanjurjo for illuminating discussions and comments on the manuscript. We would also like to thank all the participants of the workshop “Timelike Boundaries in Classical and Quantum Gravity” for stimulating discussions and the Simons Center for Geometry and Physics at Stony Brook University for its hospitality, during which this work was finalized. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities. The work of DAG is funded by UKRI Stephen Hawking Fellowship EP/W005530/1 “Quantum Emergence of an Expanding Universe”. DAG is further supported by STFC consolidated grant ST/X000753/1. RCM is also supported in part by a Discovery Grant from the Natural Sciences and Engineering Research Council of Canada, and by funding from the BMO Financial Group.
@@ A.1 Variation of the Einstein-Hilbert Action
The Einstein-Hilbert (EH) action in Euclidean signature takes the form: [^26] As usual, there is an overall minus sign in all the action expressions that follow if one wishes to work in Lorentzian signature.
$\mathcal{I}_{\text{EH}}=-\frac{1}{16\pi G_{N}}\int_{\mathcal{M}}d^{d+1}x\sqrt{g}\left(-2\Lambda+R\right)\ .$ (93)
Varying this action with respect to the inverse metric $g^{\mu\nu}$ we obtain [^27] In general, the action variation also contains a corner contribution uncovered by Hayward Hayward:1993my , whose coefficient was shown to vanish for conformal boundary conditions in Odak:2021axr . We are going to ignore corner terms in what follows, since they are irrelevant for the Euclidean computations we are after in this work.
$\delta\mathcal{I}_{\text{EH}}\hat{=}-\frac{\epsilon}{16\pi G_{N}}\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\left(\mathcal{K}_{\mu\nu}-\mathcal{K}h_{\mu\nu}\right)\delta h^{\mu\nu}-2\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{K}\right)\right)\ ,$ (94)
where $\epsilon=1(-1)$ if the boundary is timelike (spacelike) and the overhats on equal signs denote on-shell variations, as usual. Therefore, in order to have a well-posed variational principle, we have to impose boundary conditions, as well as add boundary counterterms in the action. In the following, we examine the cases of interest. As is standard in such manipulations, the following relations will be useful
$\begin{cases}\delta{h}_{\mu\nu}=-h_{\mu\sigma}h_{\nu\rho}\delta h^{\rho\sigma}\ ,\\ \delta\sqrt{h}=-\frac{1}{2}\sqrt{h}h_{\mu\nu}\delta h^{\mu\nu}\ .\end{cases}$ (95)
@@
The natural choice is to fix the induced metric along the boundary, i.e.,  impose Dirichlet boundary conditions $\delta h_{\mu\nu}|_{\partial\mathcal{M}}=0$. This kills the first term in the on-shell variation (94) and further instructs us to supplement the bulk EH action with the following Gibbons-Hawking-York (GHY) term York:1972sj ; Gibbons:1976ue :
$\mathcal{I}_{\text{GHY}}=-\frac{\epsilon}{8\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{K}\ .$ (96)
@@
Another possibility would be to fix the conjugate momentum on the boundary, which in ADM terms reads:
$\pi^{\mu\nu}:=\frac{\delta}{\delta h_{\mu\nu}}\left(\mathcal{I}_{\text{EH}}+\mathcal{I}_{\text{GHY}}\right)=-\frac{\epsilon\sqrt{h}}{16\pi G_{N}}\left(\mathcal{K}^{\mu\nu}-\mathcal{K}h^{\mu\nu}\right)\ .$ (97)
We can then massage the first term of eq. (94) as follows:
$\frac{\epsilon}{16\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\left(\mathcal{K}_{\mu\nu}-\mathcal{K}h_{\mu\nu}\right)\delta h^{\mu\nu}=-\frac{\epsilon\left(1-d\right)}{16\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}K\right)-\int_{\partial\mathcal{M}}d^{d}y\ h_{\mu\nu}\delta\pi^{\mu\nu}\ ,$ (98)
and substituting back into eq. (94):
$\delta\mathcal{I}_{\text{EH}}\hat{=}-\frac{\epsilon\left(d-3\right)}{16\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{K}\right)+\int_{\partial\mathcal{M}}d^{d}y\ h_{\mu\nu}\delta\pi^{\mu\nu}\ ,$ (99)
where we observe that subject to fixing $\delta\pi^{\mu\nu}|_{\partial\mathcal{M}}=0$ the boundary term of the EH action becomes:
$\mathcal{I}_{\text{NBCs}}^{(\text{GR})}=\frac{\epsilon\left(d-3\right)}{16\pi G_{N}}\int_{\partial\mathcal{M}}d^{d-1}y\sqrt{h}\mathcal{K}\ .$ (100)
Unlike the previous cases, the coefficient of the GHY term now depends on the dimension and, interestingly, in $d=4$ vanishes altogether, while in $d=2$ coincides with the counterterm of the Dirichlet problem.
@@
In the case of CBC (3), one fixes the conformal class of the boundary metric and the trace of the extrinsic curvature; that is, one imposes $\delta h^{\mu\nu}|_{\partial\mathcal{M}}=\Omega^{2}(y)h^{\mu\nu}$ where $\Omega(y)$ is a boundary Weyl factor and fixes $\mathcal{K}$, respectively. It is calculation wise more useful, though, to work with $\tilde{h}^{\mu\nu}:=h^{-\frac{1}{d}}h_{\mu\nu}$ such that the requirement of fixing the conformal class reduces to $\delta\tilde{h}^{\mu\nu}|_{\partial\mathcal{M}}=0$. Under a Weyl rescaling,
@@
$\tilde{h}_{\mu\nu}$ remains invariant. These boundary conditions can thus be equivalently phrased as fixing $\tilde{h}_{\mu\nu}|_{\partial\mathcal{M}},\mathcal{K}|_{\partial\mathcal{M}}$. [^28] Note that in this formulation we do not specify additional data than before, since $\tilde{h}_{\mu\nu}$ is required to have unit determinant by definition. The variations of $h_{\mu\nu}$ and $\tilde{h}_{\mu\nu}$ are related via
$\delta h_{\mu\nu}=h^{\frac{1}{d}}\delta\tilde{h}_{\mu\nu}+\frac{2}{d\sqrt{h}}h_{\mu\nu}\delta\sqrt{h}\ .$ (102)
Hence, the first term of eq. (94) takes the form:
$\int_{\partial\mathcal{M}}d^{d}y\ \pi^{\mu\nu}\delta h_{\mu\nu}$ $=\int_{\partial\mathcal{M}}d^{d}y\ h^{\frac{1}{d}}\pi^{\mu\nu}\delta\tilde{h}_{\mu\nu}+\frac{1}{d}\int_{\partial\mathcal{M}}d^{d}y\ \pi^{\mu}_{\mu}h^{\rho\sigma}\delta h_{\rho\sigma}$ (103)
$=\int_{\partial\mathcal{M}}d^{d}y\ h^{\frac{1}{d}}\pi^{\mu\nu}\delta\tilde{h}_{\mu\nu}+\frac{\epsilon(d-1)}{d8\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\ \delta\sqrt{h}\mathcal{K}\ ,$
and by virtue of a functional partial integration in field space, it becomes:
$\int_{\partial\mathcal{M}}d^{d}y\ \tilde{\pi}^{\mu\nu}\delta\tilde{h}_{\mu\nu}+\int_{\partial\mathcal{M}}d^{d}y\ \pi_{\mathcal{K}}\delta\mathcal{K}+\frac{\epsilon(d-1)}{d8\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{K}\right)\ ,$ (104)
where we defined the canonical conjugate momenta:
$\begin{cases}\tilde{\pi}^{\mu\nu}:=-\dfrac{\epsilon\sqrt{h}}{16\pi G_{N}}h^{\frac{1}{d}}\left(\mathcal{K}^{\mu\nu}-\frac{1}{d}h^{\mu\nu}\mathcal{K}\right)\ ,\\ \pi_{\mathcal{K}}:=-\dfrac{\epsilon(d-1)}{d8\pi G_{N}}\sqrt{h}\ ,\end{cases}\ ,$ (105)
and in the first line we made the conjugate momentum traceless by adding a term proportional to $h^{\mu\nu}\delta\tilde{h}_{\mu\nu}$, which is zero since:
@@
In other words, $\tilde{\pi}^{\mu\nu}$ is contracted with $\delta\tilde{h}_{\mu\nu}$ which projects to the traceless part of the corresponding tensor. Inserting (104) back in the variation (94) and imposing conformal boundary conditions, we have:
$\delta\mathcal{I}_{\text{EH}}\hat{=}\frac{\epsilon}{8d\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{K}\right)\ .$ (107)
Consequently, the variation of the bulk EH action has to be compensated by the variation of the following boundary term:
$\mathcal{I}_{\text{CBC}}=-\frac{\epsilon}{8d\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{K}\ ,$ (108)
@@
In ref. Liu:2024ymn the authors derived the necessary boundary conditions, so that the EH action has a well-posed variational problem when it is accompanied by a GHY term with an arbitrary numerical pre-factor $\Theta$, i.e.,
$\mathcal{I}_{\text{GCBC}}^{(\text{GR})}=-\frac{\Theta}{8\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{K}\ .$ (109)
@@
$\delta\mathcal{I}_{\text{EH}}\hat{=}-\frac{\epsilon}{16\pi G_{N}}\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\left(\mathcal{K}_{\mu\nu}-\mathcal{K}h_{\mu\nu}\right)\delta h^{\mu\nu}-2(1-\Theta)\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{K}\right)\right)\ ,$ (110)
and expanding the variational derivations
$\delta\mathcal{I}_{\text{EH}}\hat{=}\frac{\epsilon}{16\pi G_{N}}\left(\int_{\partial\mathcal{M}}d^{d}y\,\tilde{\pi}^{\mu\nu}\delta\tilde{h}_{\mu\nu}+2\left(1-\Theta-\frac{d-1}{d}\right)\mathcal{K}\delta\sqrt{h}+2(1-\Theta)\delta\mathcal{K}\right)\ .$ (111)
@@
$\frac{1-\Theta}{8\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}h^{-p}\delta(h^{p}\mathcal{K})\ ,$ (112)
@@
for $\Theta=1+\frac{d-1}{d(2p-1)}$. One recovers Dirichlet and conformal boundary conditions, respectively, in the limits $p\rightarrow\infty$ and $p\rightarrow 0$. In Liu:2024ymn GBCs were shown to be elliptic in Euclidean signature, for any finite $p$. Note, lastly, that GBCs have a traceless momentum (stress tensor) $\tilde{\pi}^{\mu\nu}$ canonically conjugate to the conformal class $\tilde{h}_{\mu\nu}$, for any finite $p$.

### v2

@@ Acknowledgements
It is a pleasure to thank Michael Anderson, Dionysios Anninos, Pau Figueras, Davide Gaiotto, Chawakorn Maneerat, Edgar Shaghoulian and Eva Silverstein for useful discussions. TZ especially thanks Sergio Sanjurjo for illuminating discussions and comments on the manuscript. We would also like to thank all the participants of the workshop Timelike Boundaries in Classical and Quantum Gravity for stimulating discussions and the Simons Center for Geometry and Physics at Stony Brook University for its hospitality, during which this work was finalized. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities. The work of DAG is funded by UKRI Stephen Hawking Fellowship EP/W005530/1 “Quantum Emergence of an Expanding Universe”. DAG is further supported by STFC consolidated grant ST/X000753/1. RCM is also supported in part by a Discovery Grant from the Natural Sciences and Engineering Research Council of Canada, and by funding from the BMO Financial Group.
@@ A.1 Variation of the Einstein-Hilbert Action
The Einstein-Hilbert (EH) action in Lorentzian signature takes the form [^27] We work in Lorentzian signature, so that we can have general formulas both for timelike and spacelike hypersurfaces. As usual, there is an overall minus sign in all the action expressions that follow if one wishes to work in Euclidean signature.
$\mathcal{I}_{\text{EH}}=\frac{1}{16\pi G_{N}}\int_{\mathcal{M}}d^{d+1}x\sqrt{-g}\left(-2\Lambda+R\right)\ .$ (93)
Varying this action with respect to the inverse metric $g^{\mu\nu}$ we obtain [^28] In general, the action variation also contains a corner contribution uncovered by Hayward Hayward (1993), whose coefficient was shown to vanish for conformal boundary conditions in Odak and Speziale (2021). We are going to ignore corner terms in what follows, since they are irrelevant for the Euclidean computations we are after in this work.
$\delta\mathcal{I}_{\text{EH}}\hat{=}\frac{\epsilon}{16\pi G_{N}}\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\left(\mathcal{K}_{\mu\nu}-\mathcal{K}h_{\mu\nu}\right)\delta h^{\mu\nu}-2\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{K}\right)\right)\ ,$ (94)
where $\epsilon=1(-1)$ if the boundary is timelike (spacelike) and the overhats on equal signs denote on-shell variations, as usual. Therefore, in order to have a well-posed variational principle, we have to impose boundary conditions, and add the corresponding boundary terms in the action. In the following, we examine the cases of interest. As is standard in such manipulations, the following relations will be useful
$\begin{cases}\delta{h}_{\mu\nu}=-h_{\mu\sigma}h_{\nu\rho}\delta h^{\rho\sigma}\ ,\\ \delta\sqrt{|h|}=-\frac{1}{2}\sqrt{|h|}h_{\mu\nu}\delta h^{\mu\nu}\ .\end{cases}$ (95)
@@
The natural choice is to fix the induced metric along the boundary, i.e.,  impose Dirichlet boundary conditions $\delta h_{\mu\nu}|_{\partial\mathcal{M}}=0$. This makes the first term in the on-shell variation (94) to vanish and further instructs us to supplement the bulk EH action with the following Gibbons-Hawking-York (GHY) term York (1972); Gibbons and Hawking (1977),
$\mathcal{I}_{\text{GHY}}=\frac{\epsilon}{8\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{K}\ .$ (96)
@@
Another possibility would be to fix the conjugate momentum on the boundary, which in ADM terms reads,
$\pi^{\mu\nu}:=\frac{\delta}{\delta h_{\mu\nu}}\left(\mathcal{I}_{\text{EH}}+\mathcal{I}_{\text{GHY}}\right)=-\frac{\epsilon\sqrt{|h|}}{16\pi G_{N}}\left(\mathcal{K}^{\mu\nu}-\mathcal{K}h^{\mu\nu}\right)\ .$ (97)
We can then massage the first term of eq. (94) as follows,
$\frac{\epsilon}{16\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\left(\mathcal{K}_{\mu\nu}-\mathcal{K}h_{\mu\nu}\right)\delta h^{\mu\nu}=-\frac{\epsilon\left(1-d\right)}{16\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}K\right)-\int_{\partial\mathcal{M}}d^{d}y\ h_{\mu\nu}\delta\pi^{\mu\nu}\ ,$ (98)
and substituting back into eq. (94) we obtain,
$\delta\mathcal{I}_{\text{EH}}\hat{=}\frac{\epsilon\left(d-3\right)}{16\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{K}\right)-\int_{\partial\mathcal{M}}d^{d}y\ h_{\mu\nu}\delta\pi^{\mu\nu}\ ,$ (99)
where we observe that subject to fixing $\delta\pi^{\mu\nu}|_{\partial\mathcal{M}}=0$ the boundary term of the EH action becomes
$\mathcal{I}_{\text{NBCs}}^{(\text{GR})}=-\frac{\epsilon\left(d-3\right)}{16\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{K}\ .$ (100)
Unlike the previous cases, the coefficient of the GHY term now depends on the dimension and, interestingly, in $d=3$ vanishes altogether, while in $d=1$ coincides with the boundary term of the Dirichlet problem.
@@
In the case of CBC (3), one fixes the conformal class of the boundary metric and the trace of the extrinsic curvature; that is, one imposes $\delta h^{\mu\nu}|_{\partial\mathcal{M}}=\Omega^{2}(y)h^{\mu\nu}$ where $\Omega(y)$ is a boundary Weyl factor and fixes $\mathcal{K}$, respectively. It is more useful, though, to work with $\tilde{h}^{\mu\nu}:=h^{-\frac{1}{d}}h_{\mu\nu}$ such that the requirement of fixing the conformal class reduces to $\delta\tilde{h}^{\mu\nu}|_{\partial\mathcal{M}}=0$. Under a Weyl rescaling,
@@
$\tilde{h}_{\mu\nu}$ remains invariant. These boundary conditions can thus be equivalently phrased as fixing $\tilde{h}_{\mu\nu}|_{\partial\mathcal{M}},\mathcal{K}|_{\partial\mathcal{M}}$. [^29] Note that in this formulation we do not specify additional data than before, since by definition $\tilde{h}_{\mu\nu}$ is required to have unit determinant. The variations of $h_{\mu\nu}$ and $\tilde{h}_{\mu\nu}$ are related via
$\delta h_{\mu\nu}=h^{\frac{1}{d}}\delta\tilde{h}_{\mu\nu}+\frac{2}{d\sqrt{|h|}}h_{\mu\nu}\delta\sqrt{|h|}\ .$ (102)
Hence, the first term of eq. (94) takes the form,
$\int_{\partial\mathcal{M}}d^{d}y\ \pi^{\mu\nu}\delta h_{\mu\nu}$ $=\int_{\partial\mathcal{M}}d^{d}y\ h^{\frac{1}{d}}\pi^{\mu\nu}\delta\tilde{h}_{\mu\nu}+\frac{1}{d}\int_{\partial\mathcal{M}}d^{d}y\ \pi^{\mu}_{\mu}h^{\rho\sigma}\delta h_{\rho\sigma}\,$, (103)
$=\int_{\partial\mathcal{M}}d^{d}y\ h^{\frac{1}{d}}\pi^{\mu\nu}\delta\tilde{h}_{\mu\nu}+\frac{\epsilon(d-1)}{d8\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\ \delta\sqrt{|h|}\mathcal{K}\ ,$
and by virtue of a functional partial integration in field space, it becomes,
$\int_{\partial\mathcal{M}}d^{d}y\ \tilde{\pi}^{\mu\nu}\delta\tilde{h}_{\mu\nu}+\int_{\partial\mathcal{M}}d^{d}y\ \pi_{\mathcal{K}}\delta\mathcal{K}+\frac{\epsilon(d-1)}{d8\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{K}\right)\ ,$ (104)
where we defined the canonical conjugate momenta as
$\begin{cases}\tilde{\pi}^{\mu\nu}:=-\dfrac{\epsilon\sqrt{|h|}}{16\pi G_{N}}h^{\frac{1}{d}}\left(\mathcal{K}^{\mu\nu}-\frac{1}{d}h^{\mu\nu}\mathcal{K}\right)\ ,\\ \pi_{\mathcal{K}}:=-\dfrac{\epsilon(d-1)}{d8\pi G_{N}}\sqrt{|h|}\ ,\end{cases}\ ,$ (105)
and in the first line we made the conjugate momentum traceless by adding a term proportional to $h^{\mu\nu}\delta\tilde{h}_{\mu\nu}$, which is zero since
@@
In other words, $\tilde{\pi}^{\mu\nu}$ is contracted with $\delta\tilde{h}_{\mu\nu}$ which projects to the traceless part of the corresponding tensor. Inserting (104) back in the variation (94) and imposing conformal boundary conditions, we obtain that
$\delta\mathcal{I}_{\text{EH}}\hat{=}-\frac{\epsilon}{8d\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{K}\right)\ .$ (107)
Consequently, the variation of the bulk EH action has to be compensated by the variation of the following boundary term,
$\mathcal{I}_{\text{CBC}}=\frac{\epsilon}{8d\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{K}\ ,$ (108)
@@
In ref. Liu et al. (2024) the authors derived the necessary boundary conditions, so that the EH action has a well-posed variational problem when it is accompanied by a GHY term with an arbitrary numerical pre-factor $\Theta$, i.e.,
$\mathcal{I}_{\text{GCBC}}^{(\text{GR})}=\frac{\epsilon\Theta}{8\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{K}\ .$ (109)
@@
$\delta\mathcal{I}_{\text{EH}}$ $\hat{=}\frac{\epsilon}{16\pi G_{N}}\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\left(\mathcal{K}_{\mu\nu}-\mathcal{K}h_{\mu\nu}\right)\delta h^{\mu\nu}-2(1-\Theta)\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{K}\right)\right)$ (110)
$\hat{=}-\frac{\epsilon}{16\pi G_{N}}\left(\int_{\partial\mathcal{M}}d^{d}y\,\tilde{\pi}^{\mu\nu}\delta\tilde{h}_{\mu\nu}+2\left(1-\Theta-\frac{d-1}{d}\right)\mathcal{K}\delta\sqrt{|h|}+2(1-\Theta)\sqrt{|h|}\delta\mathcal{K}\right)\ .$
@@
$\frac{1-\Theta}{8\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}h^{-p}\delta(h^{p}\mathcal{K})\ ,$ (111)
@@
for $\Theta=1+\frac{d-1}{d(2p-1)}$. One recovers Dirichlet and conformal boundary conditions, respectively, in the limits $p\rightarrow\infty$ and $p\rightarrow 0$. In Liu et al. (2024) GBCs were shown to be elliptic in Euclidean signature, for any finite $p$.